A dish of lasagna baked at 375°F is taken out of the oven into a kitchen that is 74°F. After 6 minutes, the temperature of the lasagna is 319.5°F. What will its temperature be 20 minutes after it was taken out of the oven? Round your answer to the nearest degree.
Answer by nyc_function(2741)

According to Newton's Law of Cooling, T - Ts = (To - Ts) e^(-kt), where Ts is the temperature of the surrounding medium, To is the initial temperature, T is the temperature at time t.
319.5 - 74 = (375 - 74) e^(-k • 6)
245.5 = 301 e^(-6k)
0.791935... = e^(-6k)
-0.23327... = -6k
0.038879225... = k
T - 74 = 301e^(-0.03887... • 30) = 93.7595488...
T = 167.7595488... ≈ 168°